# Triple and binary evolution pathways play equally significant roles in LISA double white dwarf formation

### v1

@@ 5.3 Possibility of electromagnetic constraints to LISA observations
While millions of double white dwarfs will emit gravitational waves in the LISA frequency bandwidth, only a few hundred double white dwarfs are well-characterized in electromagnetic observations (e.g., Munday et al., 2023). Of these, 48 of the binaries will have a high $\rho$ ratio and will serve as verification binaries for the LISA mission (e.g., Stroeer & Vecchio, 2006; Kupfer et al., 2018, 2024; Finch et al., 2023). To date, it is unknown is any of these known LISA binaries are part of triple systems with wide tertiary companion.
LISA is set to be launched in 2035 (Colpi et al., 2024). However, some electromagnetic observations/surveys planned in the near future will already increase the sample of short-period double white dwarfs and offer better constraints for the modeling, data analysis, and detection techniques of LISA double white dwarfs. This includes new binary insights from Gaia’s next data release (DR4), SDSS-V (Kervella et al., 2017), LAMOST (Zhao et al., 2012), ZTF, 4MOST (de Jong et al., 2019), WEAVE (de Vries et al., 2014), the Asteroid Terrestrial-impact Last Alert System (Heinze et al., 2018; Tonry et al., 2018), the Gravitational-wave Optical Transient Observer (Steeghs et al., 2022), Euclid (Laureijs et al., 2011), the Nancy Roman Space Telescope (Roman_2015), and Vera Rubin-LSST (Collaboration et al., 2009).
Similarly to the approach taken by Heintz et al. (2024), who used the age discrepancy from observational data to quantify the contribution of triple systems, applying comparable modeling techniques to forthcoming double white dwarf observations could provide further constraints on the contribution of triples to double white dwarfs in the LISA frequency bandwidth.
@@ 6 Conclusion
We estimated the Galactic LISA double white dwarfs from triple systems. We combined the triple population synthesis code MSE and the TNG50 cosmological simulations to seed a Milky Way-like galaxy to obtain a population of LISA-detectable double white dwarfs. To compare our results with LISA double white dwarfs from isolated binaries, we also estimated the LISA double white dwarf population from isolated binaries using the same MSE code. The main conclusions are listed below:
@@
    We identify five different key processes that shape the evolutionary pathways of triple systems, leading to the formation of double white dwarfs that emit gravitational waves in the LISA frequency bandwidth.
@@
      Triple common envelope: Mass transfer from the tertiary star onto the inner binary that eventually brings the inner binary closer to short orbital periods.
@@
      Inner binary channel: A binary with a third star that is too distant to have any effect, enters the LISA frequency bandwidth due to effectively isolated binary interactions.
@@
    Of the LISA double white dwarfs from triple systems, about $\sim 50\,\%$ of systems have a bound third star. However, the tertiary is typically too distant to have an observable imprint in the gravitational-wave signal of the inner binary.
@@
    Of the predicted LISA double white dwarfs from triple systems, we estimate that the majority of systems which enter the LISA frequency bandwidth have circular orbits, and only $3\times 10^{-6}$ (i.e. 31 systems in total in the Milky Way) of these systems have eccentric orbits. Meanwhile, in the estimated LISA double white dwarf population from isolated binary star systems, all systems have circular orbits. All the eccentric systems are found to have highly eccentric orbits ($e>0.9$) and will emit gravitational bursts, with a typical periapsis period of a few hours. However, they are unlikely to be observable due to their small gravitational-wave strain amplitude.
@@
    In our Galactic LISA double white dwarf population, we observe all types of double white dwarfs, with He-CO double white dwarfs being the dominant configuration.
Our study is the first to investigate the role of the triple evolution channel in the context of future LISA observations. We quantitatively assess the impact of including this channel on the number of observable LISA sources. While our results indicate no major distinguishable differences in the population properties of systems originating from triples compared to those formed through isolated binary evolution, they are particularly relevant for the future interpretation of LISA data on the Galactic population. Additionally, we show that the triple channel produces highly eccentric sources—although rare, these systems generate burst-like signals, in contrast to the predominantly monochromatic continuous signals emitted by the majority of the Galactic population. Thus, it is also important to consider these systems in the context of LISA data analysis.
@@ Acknowledgements.
ASR would like to thank Adrian Hamers for his guidance on the fundamentals of triple dynamics and the development of the MSE code, which was instrumental in this work. ASR also thanks Nikolaos Karnesis for constructing the confusion background using our population. JS, SJ, and SDM acknowledge funding from the Netherlands Organisation for Scientific Research (NWO), as part of the Vidi research program BinWaves (project number 639.042.728, PI: de Mink). ST also acknowledges support from the Netherlands Research Council NWO (VENI 639.041.645 and VIDI 203.061 grants). Software: MSE (Hamers et al., 2021), Matplotlib (Hunter, 2007), NumPy (van der Walt et al., 2011), pandas (McKinney, 2010), gwg (Karnesis et al., 2021). Data availability: All the codes used in this projects are avaiable at GitHub repository. The data produced during this study is available upon request.
@@ Appendix A LISA double white dwarfs from isolated binary population
In the first step, we construct a synthetic population of isolated zero-age-main-sequence binaries in the first step. The primary mass ($m_{1}$) is sampled from Kroupa’s IMF (Kroupa, 2001). Following this, we sample the orbital period of these binaries following empirically derived functions from Moe & Di Stefano (2017). Using $m_{1}$ and $T_{1}$, we sample the initial mass ratio and eccentricities using functions from Moe & Di Stefano (2017). This mass ratio is then used to calculate $m_{2}$. For a fair comparison to the triple systems, all masses are restricted to be between $1-8\,\mathrm{M_{\odot}}$. Furthermore, we reject any systems that are Roche lobe filling (Eggleton, 1983). We note that this population inherently differs from the inner binaries of triple systems. The crucial, though not the only, distinction arises from the dynamical stability requirement in triples, which forces the inner binary to be more compact. As a result, the semi-major axis distribution of inner binaries in triples is significantly more compact compared to that of isolated binaries. We repeat the process described in Sec 2.3 to create a population of $10^{5}$ isolated binaries. Based on Figure 1 we note that this isolated binary population looks distinctly different from the inner binary population of the triple systems. We evolve this synthetic population until Hubble time using MSE and select the double white dwarf that enters LISA frequency bandwidth.
In the second step, We use the same galaxy and a similar method as mentioned in Section 2.3 to seed the LISA double white dwarfs in the Galaxy. However, the total stellar mass in the simulated population is calculated after using an isolated binary population and is given by:
@@
where $N_{b,\,\mathrm{in\,range}}=10^{5}$ is the number of simulated isolated binary-star systems with MSE, $f_{b,\,\mathrm{in\,range}}$ is the fraction of binaries in this range relative to the full mass range, and $f_{t}=0.2$, $f_{b}=0.3$, and $0.51-f_{t}-f_{b}=$ 0.5 represent the fractions of triples, binaries, and singles in the full stellar population; $m_{t}=3.5\,\mathrm{M_{\odot}}$, $m_{b}=0.9\,\mathrm{M_{\odot}}$, and $m_{s}=0.5\,\mathrm{M_{\odot}}$ denote the average masses of triple, binary, and single systems, respectively.

### v2

@@ 5.3 Possibility of electromagnetic constraints to LISA observations
While millions of double white dwarfs will emit gravitational waves in the LISA frequency bandwidth, only a few hundred double white dwarfs are well-characterized in electromagnetic observations (e.g., 2023MNRAS.518.5123M). Of these, 48 of the binaries will have a high $\rho$ ratio and will serve as verification binaries for the LISA mission (e.g., 2006CQGra..23S.809S; 2018MNRAS.480..302K; 2024ApJ...963..100K; 2023MNRAS.522.5358F). To date, it is unknown is any of these known LISA binaries are part of triple systems with wide tertiary companion.
LISA is set to be launched in 2035 (LISARedBook). However, some electromagnetic observations/surveys planned in the near future will already increase the sample of short-period double white dwarfs and offer better constraints for the modeling, data analysis, and detection techniques of LISA double white dwarfs. This includes new binary insights from Gaia’s next data release (DR4), SDSS-V (2017A&A...598L...7K), LAMOST (2012arXiv1206.3569Z), ZTF (2019PASP..131a8002B) , 4MOST (2019Msngr.175....3D), WEAVE (2014MNRAS.438.1909D), the Asteroid Terrestrial-impact Last Alert System (2018AJ....156..241H; 2018PASP..130f4505T), the Gravitational-wave Optical Transient Observer (2022MNRAS.511.2405S), Euclid (Euclid_2011), the Nancy Roman Space Telescope (Roman_2019), and Vera Rubin-LSST (LSST_2009). Similarly to the approach taken by 2024ApJ...969...68H, who used the age discrepancy from observational data to quantify the contribution of triple systems, applying comparable modeling techniques to forthcoming double white dwarf observations could provide further constraints on the contribution of triples to double white dwarfs in the LISA frequency bandwidth.
@@ 6 Conclusion
We estimated the Galactic LISA double white dwarfs from triple systems. We combined the triple population synthesis code MSE and the TNG50 cosmological simulations to seed a Milky Way-like galaxy and obtain a population of LISA-detectable double white dwarfs. To compare our results with LISA double white dwarfs from isolated binaries, we also estimated the LISA double white dwarf population from isolated binaries using the same MSE code. Our main conclusions are listed below:
@@
    We identify five different key processes that shape the evolutionary pathways of triple systems and lead to the formation of double white dwarfs that emit gravitational waves in the LISA frequency bandwidth.
@@
      Triple common envelope: The tertiary star overflows its Roche lobe and transfers mass onto the inner binary, triggering a triple common envelope phase that tightens the inner orbit to short periods.
@@
      Inner binary channel: A binary with a third star that is too distant to have any effect enters the LISA frequency bandwidth due to effectively isolated binary interactions.
@@
    Of the LISA double white dwarfs from triple systems, about $50\,\%$ of the systems have a bound third star. However, the tertiary is typically too distant to have an observable imprint in the gravitational wave signal of the inner binary.
@@
    Of the predicted LISA double white dwarfs from triple systems, we estimate that the majority of the systems that enter the LISA frequency bandwidth have circular orbits, and only $3\times 10^{-6}$ (i.e., 31 systems in total in the Milky Way) of these systems have eccentric orbits. Meanwhile, in the estimated LISA double white dwarf population from isolated binary star systems, all systems have circular orbits. All the eccentric systems are found to have highly eccentric orbits ($e>0.9$) and will emit gravitational bursts, with a typical periapsis period of a few hours. However, they are unlikely to be observable due to their small gravitational wave strain amplitude.
@@
    In our Galactic LISA double white dwarf population, we observe all types of double white dwarfs. Nonetheless, containing a He white dwarf and a CO white dwarf is the most common configuration.
Our study is the first to investigate the role of the triple evolution channel in the context of future LISA observations. We quantitatively assessed the impact of including this channel on the number of observable LISA sources. While our results indicate no major distinguishable differences in the population properties of systems originating from triples compared to those formed through isolated binary evolution, they are particularly relevant for the future interpretation of LISA data on the Galactic population. Additionally, we show that the triple channel produces highly eccentric sources; although rare, these systems generate burst-like signals, in contrast to the predominantly monochromatic continuous signals emitted by the majority of the Galactic population. Thus, it is also important to consider these systems in the context of LISA data analysis.
Data availability
The catalog used to generate the astrophysical noise background from double white dwarfs (both triple and isolated binary origin) is available at Zenodo.
@@ Acknowledgements.
ASR would like to thank Adrian Hamers for his guidance on the fundamentals of triple dynamics and the development of the MSE code, which was instrumental in this work. ASR also thanks Nikolaos Karnesis for constructing the confusion background using our population. We thank Onno Pols for clarifying to us the assumptions adopted in the binding energy parameter ($\lambda$) fits provided by 2014A&A...563A..83C. JS, SJ, and SDM acknowledge funding from the Netherlands Organisation for Scientific Research (NWO), as part of the Vidi research program BinWaves (project number 639.042.728, PI: de Mink). ST also acknowledges support from the Netherlands Research Council NWO (VENI 639.041.645 and VIDI 203.061 grants). Software: MSE (2021MNRAS.502.4479H), Matplotlib (2007CSE.....9...90H), NumPy (2011CSE....13b..22V), pandas (mckinney2010), gwg (kar21).
@@ Appendix A LISA double white dwarfs from isolated binary population
In the first step, we construct a synthetic population of isolated zero-age main-sequence binaries. The primary mass ($m_{1}$) is sampled from Kroupa’s IMF (2001MNRAS.322..231K). Following this, we sample the orbital period of these binaries following empirically derived functions from 2017ApJS..230...15M. Using $m_{1}$ and $T_{1}$, we sample the initial mass ratio and eccentricities using functions from 2017ApJS..230...15M. This mass ratio is then used to calculate $m_{2}$. For a fair comparison to the triple systems, all masses are restricted to be between $1-8\,\mathrm{M_{\odot}}$. Furthermore, we reject any systems that are Roche lobe filling (1983ApJ...268..368E). We note that this population inherently differs from the inner binaries of triple systems. The crucial, though not the only, distinction arises from the dynamical stability requirement in triples, which forces the inner binary to be more compact. As a result, the semi-major axis distribution of inner binaries in triples is significantly more compact compared to that of isolated binaries (2023ApJ...950....9R). We repeat the process described in Sect. 2.3 to create a population of $10^{5}$ isolated binaries. Based on Fig. 1 we note that this isolated binary population looks distinctly different from the inner binary population of the triple systems. We evolve this synthetic population until Hubble time using MSE and select the double white dwarf that enters LISA frequency bandwidth.
In the second step, We use the same galaxy and a similar method as mentioned in Sect. 2.3 to seed the LISA double white dwarfs in the Galaxy. However, the total stellar mass in the simulated population is calculated after using an isolated binary population and is given by
@@
where $N_{b,\,\mathrm{in\,range}}=10^{5}$ is the number of simulated isolated binary-star systems with MSE, $f_{b,\,\mathrm{in\,range}}$ is the fraction of binaries in this range relative to the full mass range, and $f_{t}=0.2$, $f_{b}=0.3$, and $1-f_{t}-f_{b}=0.5$ represent the fractions of triples, binaries, and singles in the full stellar population; $m_{t}=3.5\,\mathrm{M_{\odot}}$, $m_{b}=0.9\,\mathrm{M_{\odot}}$, and $m_{s}=0.5\,\mathrm{M_{\odot}}$ denote the average masses of triple, binary, and single systems, respectively. This gives an estimate of about $\sim 3.8\times 10^{6}$ LISA double white dwarfs.
To assess the impact of the triple fraction on normalization, we repeat the same procedure with zero triple fraction ($f_{t}=0$). Under this assumption, the estimated number of Galactic LISA double white dwarfs originating from isolated binaries is approximately $\sim 9\times 10^{6}$. All these results are presented in Table 1.
Appendix B Estimation of sampling uncertainty
We estimate the sampling uncertainty introduced by the stochastic seeding of the galaxy in our population using the bootstrapping method. Specifically, we repeat the procedure described in Sect. 2.3 fifty times. In each iteration, we sample from the intrinsic MSE double white dwarf population, separately for systems with triple and isolated binary origins, with replacement (replacement=True). This approach allows for random selection with repeated entries from each population. As a result, we generate fifty distinct mock populations for Galactic double white dwarfs from systems with triple origins and fifty more from systems with isolated binary origins.
To quantify the uncertainty, we calculate error bars based on the $\approx 15.9$th and $\approx 84.1$th percentiles of the mock populations, corresponding to the $1\sigma$ ($\approx 68.3\%$) confidence interval.